OpenStudy (anonymous):

Triangle DEF has sides with lengths 8, 9, and 11. Is the triangle obtuse, acute, or right?

OpenStudy (anonymous):

@agentx5

OpenStudy (anonymous):

Remember this? \[\frac{a}{\sin \angle A}+\frac{b}{\sin \angle B}=\frac{c}{\sin \angle C}\]  Then you can use Heron's formula to find a base & height, the base will be one of the sides. Area\(_{\triangle}\)=\(\large\sqrt{s(s-a)(s-b)(s-c)}\) where \(\large s= \frac{a+b+c}{2}\) And area is also = \(\frac{1}{2}(base)(height)\)